The density of ethanol, a colorless liquid that is commonly known as alcohol, is 0.798 g/mL. Calculate the mass of 17.4 mL of ethanol.
1
Expert's answer
2017-08-30T09:53:07-0400
The mass is derived as a density multiplied by a volume. Thus, m = d*V = 0.798 g/mL*17.4 mL = 13.9 g